OpenStudy (anonymous):

can someone please help me simplify x^2-2/4x+1/16

OpenStudy (anonymous):

is the middle term (2/4)x ?

OpenStudy (anonymous):

you need to recognized that this expression can be factored... key "clues" include the x^2 term and the (1/16) term...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(a-b)²=a²-2ab+b²

OpenStudy (anonymous):

x^2-(2/4x)+(1/16)

OpenStudy (anonymous):

Do you see how the form of @RaphaelFilgueiras's expression above is the same form as your problem?

OpenStudy (anonymous):

b^2 in his expression is (1/16) in yours... especially if you realize that (1/4)^2 is = (1/16)

OpenStudy (anonymous):

so 1/16 is simplified to (1/4)^2. what happens to 2/4x

OpenStudy (anonymous):

it can be reduced to 1/2, i know that...

OpenStudy (anonymous):

It isn't simplified... you just need to write it in that format to make it "something squared". See how @RaphaelFilgueiras listed the equation with a b^2 at the end? You needed your problem to have "something squared" at the end... you started with (1/16), but it becomes "something squared" if you rewrite it as (1/4)^2

OpenStudy (anonymous):

the middle term is 2ab, nothing squared

OpenStudy (anonymous):



OpenStudy (anonymous):

isnt that only the 1st and 3rd term

OpenStudy (anonymous):



OpenStudy (anonymous):

yes, but actually don't simplify to (1/2)... just leave it as 2/4. (a-b)(a-b) = a^2 - ab - ab + b^2 = a^2 - 2ab + b^2 In your case, you have a middle term of -(2/4)x. -(2/4)x would be exactly like -2ab if a = x and b = 1/4 Try multiplying out (x - 1/4)(x - 1/4) and you'll see it matches your problem :)

OpenStudy (anonymous):

i see..

OpenStudy (anonymous):

thanks